Phillip charged $400 worth of goods on his credit card.on his first bill, he was not charged any interest, and he a payment of $20. He then charged $18 worth of goods. On his second bill a month later, he was charged 2% interest on his entire unpaid balance. How much interest was Philip charged on his second bill?
Answer by TimothyLamb(4379)

first month balance carried over:
b = 400 - 20
b = 380
---
second month billed:
b = 380 + 18
b = 398
---
second month interest charged:
i = 398*0.02
i = 7.96
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---